Mothballs are composed primarily of the hydrocarbon naphthalene (C10H8). When 1.229 g
g of naphthalene burns in a bomb calorimeter, the temperature rises from 24.604 ∘C to 30.460 ∘C.

Find ΔrU for the combustion of naphthalene. The heat capacity of the calorimeter, determined in separate experiment, is 5.112 kJ∘C−1.
1
Expert's answer
2020-12-18T08:46:01-0500

The heat produced when the naphthalene burns can be calculated from the heat capacity of the calorimeter, 5.112 kJ/°C and the temperature change (30.460-24.604)°C:

"Q = c\u2206T = 5.112\u00b7(30.460-24.604) = 29.94" kJ.

The bomb calorimetry condition is constant volume condition. Therefore, the heat released corresponds to the internal energy change "\u2206U" . To calculate "\u2206_rU" , the heat must be divided by the number of the moles of the naphthalene heated.

The number of the moles "n" of naphthalene is its mass divided by its molar mass "M = 128.17" g/mol:

"n = \frac{m}{M} = \frac{1.229}{128.17}=0.009589" mol.

Finally, the internal energy change of the reaction is (the negative sign indicates that the reaction is exothermic):

"\Delta_rU =- \frac{Q}{n} =- \frac{29.94}{0.009589} =- 3122" kJ/mol.


Answer: "\Delta_rU" for the combustion of naphthalene is -3122 kJ/mol.